40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 195 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  7352   Thu Sep 6 17:11:40 2012 janosch, Manasa, SteveUpdate pick-off and baffle at ETMY

We have installed the pick-off mirror at the ETMY table for the small-angle scattering measurement on ITMY. As we had already done for the X arm pick-off, the pick-off mirror at ETMY was aligned shooting a green laser normally through the viewport on the pick-off and steering it onto ITMY.

A baffle was also installed at a distance of about 30cm from ETMY near the edge of the table.

  7351   Thu Sep 6 17:06:25 2012 Rijuparna ChakrabortyConfigurationelogCavitymode scan

 Aim: to scan the cavitymodes of IMC

The circuit used: 

 Attachment4

Results obtained:

Attachment 1,2,3

 

Attachment 1: 3.pdf
3.pdf
Attachment 2: 2.pdf
2.pdf
Attachment 3: 1.pdf
1.pdf
Attachment 4: cavityscanconnections.pdf
cavityscanconnections.pdf
  7350   Thu Sep 6 16:46:44 2012 JenneUpdateSUSBS aligned, target removed

Q and I aligned the BS such that we were hitting the center of ETMX. The ETMX cage does not have OSEM setscrew holes on the front, so it is not possible to put the targets that Steve made on this optic.  So, I put the freestanding ruler in front of the optic, with the edge of the ruler at the center (as viewed from above) of the optic.  Then Eric steered the BS until we were hitting the 5.5" mark, and roughly half of the beam was obscured by the ruler.

We then aligned ITMX such that the prompt reflection was colinear with the incoming beam. 

I checked the 2 spots through the BS, heading to the AS port.  (2 spots since MICH hasn't been locked / finely aligned yet).  They were being clipped on the 2nd output PZT.  I adjusted the knobs of the first output PZT to center the spots on the 2nd PZT.  Note that the output PZTs' power is still off, and has been off for some unknown length of time.  I had found them off when prepping for the vent a week or two ago.  So the current alignment depends on them staying off.  We don't really need them on until we're ready to employ our OMC.

The beams now look nicely unclipped on the AS camera, and we're aligning MICH.

  7349   Thu Sep 6 13:07:02 2012 JenneUpdateIOOIPANG no longer a reference :(

I was having trouble centering IPANG using the PZTs, and I suspected something funny was going on at the end.  I went down there, and the beam was focused right on the PD, and the spot was very very small.  I think this means that when I was trying to center the beam, I was falling into the gap between the pieces of the diode.  Also, as Koji pointed out to me the other day, if the PD is at the focal point of the beam, any parallel rays hitting the lens just before the PD will all go to the same place, no matter how the input beam has moved.  This means we're not getting as much info out as we'd like.

So.  I moved the lens a little bit farther from the PD such that we are just beyond the focal point of the beam.  The beam size is now ~1mm on the QPD.

This means, however, that I moved the beam on the QPD such that IPANG is no longer a reference of the input pointing. Ooops. I think this adjustment needed to be done though.  Right now, the PZTs are set to where we had them yesterday, when we moved them slightly to center the IPANG QPD, and I've recentered IPANG.

  7348   Thu Sep 6 10:57:27 2012 JenneUpdateGeneralForgot to turn green refl pd back on

Quote:

I couldn't understand the Y-End green setup as the PD was turned off and the sign of the servo was flipped. Once they are fixed, I could lock the cavity with the green beams.

Quote:

[EricQ, Jenne, brains of other people]

Get green spots co-located with IR spots on ETMs, ITMs, check path of leakage through the arms, make sure both greens get out to PSL table

 

 I had turned the green refl PD off on Tuesday while we were doing the IPANG alignment, since the beam was not so bright, and the LED on top of the PD was very annoyingly bright.  I forgot to turn it back on.  The sign flip on the servo, I can't explain.

  7347   Thu Sep 6 02:51:47 2012 KojiUpdateGeneralGreen beam roughtly aligned (Re: Yarm aligned to IR incident beam)

The Y-End green beam was roughly aligned by the steering mirrors for the green beam.

I couldn't understand the Y-End green setup as the PD was turned off and the sign of the servo was flipped. Once they are fixed, I could lock the cavity with the green beams.

After a long alignment session, TEM00 was found. The alignment of the green beam has not been optimized.

Looking at the spot position at ETMY OSEM holders (not by the ccd image), it seems that the cavity mode is not at the center of the mirrors.

Quote:

[EricQ, Jenne, brains of other people]

Get green spots co-located with IR spots on ETMs, ITMs, check path of leakage through the arms, make sure both greens get out to PSL table

 

Attachment 1: green_y.png
green_y.png
  7346   Wed Sep 5 19:29:45 2012 JenneUpdateGeneralYarm aligned to IR incident beam

[EricQ, Jenne, brains of other people]

Checked at ETMY that the pointing hadn't changed a whole lot.  Jamie and Koji pointed out that we were half falling off of the IPANG QPD.  Adjusted PZT2 sliders so that we were again centered on IPANG's QPD.  Before we close up, we'll want to put the sliders back to their nominal positions, and use the knobs to hit IPANG, but this is equivalent and fine for now.  The tip tilts don't seem to have moved much overnight, since the beam drift on both IPANG and ETMY was fixed simultaneously with PZT2 (recall, IPANG picked off before tip tilts exist in the beam path).  This left us hitting the center of ETMY.  We moved ETMY sliders to make the reflected beam hit the center of ITMY (same spot position as transmitted beam from BS).  Then moved ITMY to get prompt reflection to hit same spot on ETMY as original primary beam from BS.  Checked at ITMY that we didn't need to move ETMY anymore.  (Actually, I forget how many iterations we did, but in the end, all of the reflections that we can find are co-located on the test masses.)

Next up:

Align BS so we're hitting the center of ETMX

Tap / readjust ITMX OSEM which is at 0.3 to get it back to the center of its range

Align ITMX to lock MICH

Check no clipping on POX / POY optics, no clipping around BS

Check PRM, SRM alignment (what exactly do we want to do here? Try to lock PRMI / SRMI?)

Get IPPOS out of vac

Fix clipped ITMY / SRM oplev

Install 'black' glass beam dumps - forward-going POP beam, 2 places in BS chamber (check old elog from Jenne/Yuta for the places).

Get green spots co-located with IR spots on ETMs, ITMs, check path of leakage through the arms, make sure both greens get out to PSL table

  7345   Wed Sep 5 13:11:43 2012 jamieOmnistructureVACbetter in-air "lite" access connector needed

Quote:

Quote:

Quote:

We really need something better to replace the access connector when we're at air.  This tin foil tunnel crap is dumb.  We can't do any locking in the evening after we've put on the light doors.  We need something that we can put in place of the access connector that allows us access to the OMC and IOO tables, while still allowing IMC locking, and can be left in place at night.

 It is in the shop. It will be ready for the next vent. Koji's dream comes through.

 Can we see the full design?  If we can't lock the mode cleaner with this thing on then it's really of no use.  We want it to be equivalent to the light doors, but allow us to keep the mode cleaner locked.  That's the most important aspect.

 It also needs to be wide enough that the MMT beam can go through, so that we can not only lock the MC, but also work on the rest of the IFO.

  7344   Wed Sep 5 10:50:15 2012 jamieOmnistructureVACbetter in-air "lite" access connector needed

Quote:

Quote:

We really need something better to replace the access connector when we're at air.  This tin foil tunnel crap is dumb.  We can't do any locking in the evening after we've put on the light doors.  We need something that we can put in place of the access connector that allows us access to the OMC and IOO tables, while still allowing IMC locking, and can be left in place at night.

 It is in the shop. It will be ready for the next vent. Koji's dream comes through.

 Can we see the full design?  If we can't lock the mode cleaner with this thing on then it's really of no use.  We want it to be equivalent to the light doors, but allow us to keep the mode cleaner locked.  That's the most important aspect.

  7343   Wed Sep 5 09:50:25 2012 SteveOmnistructureVACbetter in-air "lite" access connector needed

Quote:

We really need something better to replace the access connector when we're at air.  This tin foil tunnel crap is dumb.  We can't do any locking in the evening after we've put on the light doors.  We need something that we can put in place of the access connector that allows us access to the OMC and IOO tables, while still allowing IMC locking, and can be left in place at night.

 It is in the shop. It will be ready for the next vent. Koji's dream comes through.

Attachment 1: IMG_1612.JPG
IMG_1612.JPG
  7342   Tue Sep 4 20:25:22 2012 jamieOmnistructureVACbetter in-air "lite" access connector needed

We really need something better to replace the access connector when we're at air.  This tin foil tunnel crap is dumb.  We can't do any locking in the evening after we've put on the light doors.  We need something that we can put in place of the access connector that allows us access to the OMC and IOO tables, while still allowing IMC locking, and can be left in place at night.

  7341   Tue Sep 4 20:20:47 2012 jamieUpdateGeneralproblematic tip-tilts

Quote:

We clearly need a better plan for adjusting the tip tilts in pitch, because utilizing their hysteresis is ridiculous.  Koji and Steve are thinking up a set of options, but so far it seems as though all of those options should wait for our next "big" vent.  So for now, we have just done alignment by poking the tip tilt.

Tomorrow, we want to open up the MC doors, open up ETMY, and look to see where the beam is on the optic.  I am concerned that the hysteresis will relax over a long ( >1hour ) time scale, and we'll loose our pointing.  After that, we should touch the table enough to trip the BS, PRM optics, since Koji is concerned that perhaps the tip tilt will move in an earthquake.  Jamie mentioned that he had to poke the tip tilt a pretty reasonable amount to get it to change a noticeable amount at ETMY, so we suspect that an earthquake won't be a problem, but we will check anyway.

 I'm very unhappy with the tip-tilts right now.  The amount of hysteresis is ridiculous.  I have no confidence that they will stay pointing wherever we point them.  It's true I poked the top more than it would normally move, but I don't actually believe it wouldn't move in an earthquake.  Given how much hysteresis we're seeing, I expect it will just drift on it's own and we'll loose good pointing again.

And as a reminder, IPPOS/ANG don't help us here before the tip-tilts are in the PRC after the IP pointing sensors.

I think we need to look seriously at possible solutions to eliminate or at least reduce the hysteresis, by either adding weight, or thinner wire, or something.

  7340   Tue Sep 4 20:13:46 2012 JenneUpdateGeneralThe Plan - Tues evening version

 Tues

* Hit center of ETMY, using input optics, PR3.

* Get IPANG out of vac, center QPD.

Wed

* AM: Riju do MC mode scans

* Starting right after 40m meeting, if not before: Lock MICH to get BS, ITMs aligned well

* Check if beam is hitting center of ITMX.

* Check for clipping around BS

     - Use Watek to look at beam at all 4 BS ports - make sure no clipping going into BS, after BS in the michelson, or the AS port

       - Use some old in-vac mirrors to direct beam out the BS door.  Cameras are waiting near BS chamber.

* Prepare glass beam dumps??

* IPPOS - make sure beam gets out of chamber

* Jan take photos of ETMX scattering setup

* Manasa take in-vac photos of all tables, for table layouts

* Jan / Manasa viewport transmission

* Install glass beam dumps?

* Install glass baffle at ETMY. Jan maybe install baffle at one of ITMs.

* ????

Thurs

* ????

* Check table levelling one last time on all tables. 

Fri

* Close all heavy doors.  (Access connector, ITMX, ITMY, BS, ETMX, ETMY? )

* Drag wipe test masses

* Start at ~10am?

Mon (if not Fri)

* Start pumping

  7339   Tue Sep 4 20:06:04 2012 JenneUpdateLockingMC scan input switched to the 11MHz port of EOM

Since the EOM's signal combiner (splitter backwards) is frequency-independent, Koji and Jamie (in the proper turn off, turn on order) put the 55MHz signal back to the EOM, and put the MC mode scan input to the 11MHz port.  This way we can lock the Michelson tomorrow, and we don't have to keep switching cables around when Riju wants to take some scans.

  7338   Tue Sep 4 20:03:38 2012 JenneUpdateGeneralin-vac progress

Opened ETMY, beam was high and to the right (if you look at the face of ETMY).  Tried walking beam up, since just doing PZT2 caused clipping at the BS before we got to the correct spot on ETMY.  Moved PZT1, then PZT2, to translate beam, but we couldn't get far enough without starting to fall off of PZT2.  Put PZT1 approx. back where it was. Jamie tapped on the top of PR3 (tip tilt just before BS), and then did some compensation with PZT2 to get the beam through the BS target to the center of ETMY.

The beam is very dim at ETMY.  We ended up holding the big IR card with holes such that one of the holes was near the center of the optic, in front of the cage.  Then one person turned off all the room lights so we could see the beam, another person moved PZT2 and PR3, then lights person turned on and off the lights so we could compare beam position with hole position.  A total pain, but it ended up working better than just trying to follow the beam with a card.

We clearly need a better plan for adjusting the tip tilts in pitch, because utilizing their hysteresis is ridiculous.  Koji and Steve are thinking up a set of options, but so far it seems as though all of those options should wait for our next "big" vent.  So for now, we have just done alignment by poking the tip tilt.

Tomorrow, we want to open up the MC doors, open up ETMY, and look to see where the beam is on the optic.  I am concerned that the hysteresis will relax over a long ( >1hour ) time scale, and we'll loose our pointing.  After that, we should touch the table enough to trip the BS, PRM optics, since Koji is concerned that perhaps the tip tilt will move in an earthquake.  Jamie mentioned that he had to poke the tip tilt a pretty reasonable amount to get it to change a noticeable amount at ETMY, so we suspect that an earthquake won't be a problem, but we will check anyway.

After doing all of that, we adjusted IPANG so that the beam gets out of the vacuum envelope.  The beam must squeeze between the wall of the black beam tube and a lens for the oplev, so there is a very, very limited amount of space.  The eventual new tables will be very helpful for alleviating this, but for now we must live with it.  Even though it is pretty squishy right at the edge of the table, with the new layout we should think about giving IPANG a little more space.  Basically, if there is any Yaw motion, the beam going to the QPD will be clipped, and we may get confusing info.  We moved some of the IPANG optics that are on the end transmission table so the beam is centered on the optics while it just makes it through the space between the wall and the lens.  The spot was centered on the IPANG QPD. 

We still need to check on IPPOS, but it is always easier than IPANG.

General notes:

We discovered that the 45deg SOS beam targets are awesome, especially the ones with the irises.  The plain hole ones have very small holes relative to our beam size, so they are much more useful for the MC optics (which they were designed for).  The 45deg targets were made such that the target holes do not line up with the mounting holes.  This is good, since the mounting holes are lower than the center of the optic.  I don't think ericQ and I realized that on Friday, so it's probable that we had installed the target upside-down.  We still need to remake the 0deg targets for the next vent.

Jamie dogged down the new 'bathroom mirror' that lets us see BS and PRM on the same camera view.  He also adjusted some of the masses on the BS table to relevel the table. We need to (at some point) rename the AS_SPARE camera to something like BS_PRM, since we plugged the new camera into the AS_SPARE port on the videoswitch.

tl;dr: Input beam adjusted so we're hitting center of ETMY.  IPANG coming out of vacuum, QPD centered. Need better tip tilt in-situ adjustment capability.

  7337   Tue Sep 4 13:50:26 2012 ericqUpdatePSLPMC Realigned, power adjusted

 I adjusted the PMC alignment this morning, brought the transmission up to 0.83V.

After the lunch meeting, we found the the MC transmission was higher than recently seen. Turned out the HWP had drifted, causing 30mW to be input to the MC. I adjusted it back down to 20mW. 

  7336   Tue Sep 4 13:44:17 2012 ManasaUpdateGeneralThe Plan

Quote:

We need a plan for the rest of the week.  I want to be closing the heavy doors on Friday at the latest.  Please add to / comment on this list!

 

Tues

* Lock MICH to get BS, ITMs aligned well

* Check if beam is hitting center of ITMs. 

* Check for clipping around BS

     - Use Watek in-vac to look at beam at all 4 BS ports - make sure no clipping going into BS, after BS in the michelson, or the AS port

* Try to get arms to flash??

* Prepare glass beam dumps??

Wed

* IPPOS / IPANG - make sure beam gets out of chambers (this may require opening ETMY)

* Jan take photos of ETMX scattering setup

* Manasa take in-vac photos of all tables, for table layouts

* Install glass beam dumps?

* If ETMY open, install glass baffle

* ????

Thurs

* ????

* Check table levelling one last time on all tables. 

Fri

* Close all heavy doors.  (Access connector, ITMX, ITMY, BS, ETMX, ETMY? )

* Start at ~10am?

Mon (if not Fri)

* Start pumping

Wed

* Jan/Manasa - Measure transmission of viewport at ETMX

 

 

 

 

 

  7335   Tue Sep 4 13:31:55 2012 JenneUpdateGeneralThe Plan

We need a plan for the rest of the week.  I want to be closing the heavy doors on Friday at the latest.  Please add to / comment on this list!

 

Tues

* Lock MICH to get BS, ITMs aligned well

* Check if beam is hitting center of ITMs. 

* Check for clipping around BS

     - Use Watek in-vac to look at beam at all 4 BS ports - make sure no clipping going into BS, after BS in the michelson, or the AS port

* Try to get arms to flash??

* Prepare glass beam dumps??

Wed

* IPPOS / IPANG - make sure beam gets out of chambers (this may require opening ETMY)

* Jan take photos of ETMX scattering setup

* Manasa take in-vac photos of all tables, for table layouts

* Jan / Manasa viewport transmission

* Install glass beam dumps?

* If ETMY open, install glass baffle

* ????

Thurs

* ????

* Check table levelling one last time on all tables. 

Fri

* Close all heavy doors.  (Access connector, ITMX, ITMY, BS, ETMX, ETMY? )

* Drag wipe test masses

* Start at ~10am?

Mon (if not Fri)

* Start pumping

  7334   Tue Sep 4 11:32:58 2012 JenneUpdateLockingFriday in-vac work

Elog re: Friday's work

Adjusted PZT2 so we're hitting the center of PR2. 

Noticed that the beam centering target is too low by a few mm, since the OSEM set screw holes that it mounts to are lower than the center line of the optic.  This meant that while we were hitting the center of PR2, the beam was half clipped by PRM's centering target.  We removed the target to confirm that the beam is really centered on PR2.

Checked the beam on PR3 - it looked fine.  There had been concern last week that PR2 was severely pitched forward, but this turns out to be an effect of the PRM centering target being too low - shoot the beam downward to go through the hole, beam continues downward to hit the bottom of PR2, so beam is falling of the bottom of PR3.  But when we actually centered the beam on PR2, things looked fine on PR3. 

Checked that the beam approximately goes through the beam splitter.  Again, the targets are too low, and these 45 deg targets' holes are smaller than the 0 deg targets, so we don't see any beam going through the target, since the beam is hitting the target higher than the hole.  The beam looked left/right like it was pretty close to the hole, but it was hard to tell since the angle is bad, and I'm not infinitely tall.  We should check again to make sure that the beam is going through properly, and we're not clipping anywhere.  I'll need help from a height-advantaged person for this.

Checked that the beam is hitting the center of the ITMY, as best we can see by using an IR card at the back of the optic.  We didn't try reaching around to put a target on the front side. 

We were debating whether it would be worth it to open ETMY this week, to check that the beam transmitted through the BS hits the center of ETMY.

We also took a quick look around the AS optics, but since that depends on BS/ITMX alignment, we weren't sure how to proceed.  We need a plan for this part.  All suspended optics were restored to their last good alignment, but we haven't tried locking MICH or anything to confirm that the alignment. 

To do list:  Check no clipping on ITMY table of beam between BS and ITMY, clipping on POY optics.  Also, oplev is clipping on cable holder thing on the table - this needs to be moved.  .....other?

  7333   Tue Sep 4 10:29:41 2012 SteveUpdateSUSPRM damping restored
  7332   Fri Aug 31 18:32:47 2012 KojiUpdateIOOCorrecting improper termination of the 55MHz input to the EOM

 

The power switch of the frequency source was mechanically broken. I replace it with the spare in Kiwamu's treasure box.
I think this replaced new one also had the same problem before, we basically have to find an alternative switch.
The problem is too much friction between the metal housing and the plastic button. I have lubricated the button with
bicycle oil.

I have checked that the 55MHz output still have the output of 25dBm after a 5m BNC cable. (nominal: 27dBm)

I could not test it with the MC as the chambers have already been closed with the light doors.

  7331   Fri Aug 31 17:50:41 2012 SteveUpdateSUSSOS centering target

The SOS centering target is 1.9 mm lower than it should be! 

The hole is 10mm for the  ~6 mm beam

 

 

 

  7330   Fri Aug 31 17:44:21 2012 ManasaUpdateRingdownData

Quote:

Ok, so the whole idea that mirror motion can explain the ripples is nonsense. At least, when you think off the ringdown with "pump off". The phase shifts that I tried to estimate from longitudinal and tilt mirror motion are defined against a non-existing reference. So I guess that I have to click on the link that Koji posted...

Just to mention, for the tilt phase shift (yes, there is one, but the exact expression has two more factors in the equation I posted), it does not matter, which mirror tilts. So even for a lower bound on the ripple time, my equation was incorrect. It should have the sum over all three initial tilt angles not only the two "shooting into the long arms" of the MC.

Quote:

Laser frequency shift = longitudinal motion of the mirrors

Ringing: http://www.opticsinfobase.org/ol/abstract.cfm?uri=ol-20-24-2463

Quote:

Hmm. I don't know what ringing really is. Ok, let's assume it has to do with the pump... I don't see how the pump laser could produce these ripples. They have large amplitudes and so I always suspected something happening to the intracavity field. Therefore I was looking for effects that would change resonance conditions of the intracavity field during ringdown. Tilt motion seemed to be one explanation to me, but it may be a bit too slow (not sure yet). Longitudinal mirror motion is certainly too slow. What else could there be?

 

 

It is essential we take a look at the ringdown data for all measurements made so far to figure out what must be done to track the source of these notorious ripples. I've attached the plot for the same showing the decay time to be the same in all cases. About the ripples; it seems unlikely to both Jan and me that the ripples are some electronic noise because the ripples do not follow any common pattern or time constant. We have discussed with Koji about monitoring the frequency shift, the input power to the MC and also try other methods of shutting down the pump to track their source as the next steps.

 

cum_plot.png 

  7329   Fri Aug 31 17:26:54 2012 SteveUpdateGeneralBS camera

Steve and Eric

Placed pick up mirror for BS face and PRM back. I will ask Jamie to clamp it.

There will be an other camera set up to view the face of PRM

Attachment 1: BS_PRM.png
BS_PRM.png
  7328   Fri Aug 31 13:22:29 2012 ElliUpdateIOOCorrecting improper termination of the 55MHz input to the EOM

Koji, Riju, Elli

This morning Koji discovered that the 55MHz input into the RF combiner that I disconnected yesterday wasn't terminated properly, so it was reflecting power back into the amplifier in the signal generation unit.    We turned off the signal generation unit and checked that the amplifier was still working properly- it was.  A 50 ohm terminator was attached to the end of the 55MHz cable so that it is now terminated properly.

When we tried to turn the signal generator box back on we discovered the switch is broken (the box will only stay on while you hold down the on switch) and will need to be replaced.  In order to create the 29.5MHz sidebands to lock the mode cleaner, we bypassed the signal generation unit which won't stay on (unplugging '29.5 MHZ out' cable from the frequency generation unit), and instead sent a 0.39V 29.5MHz signal from a function generatior into 'RF input' on the 'RF AM Stabiliser' board.

We also increased the power coming exiting the PSL table and going into the cavity from 11 microwatts to 20 microwatts by adjusting the polariser at the end of the table slightly.  The power has been set to 20 microwatts using the polariser a few days ago but had drifted down since then.

  7327   Fri Aug 31 10:24:36 2012 SteveUpdateVACchamber dog clamps checked

Quote:

I tightened as many of the dog clamps on the bottom of the BS, ITMX and ITMY chambers as I could find.  I used a torque wrench at 45 ft-lbs.  Some of the bolts of the dogs were too long, and I couldn't find an extender thing to accommodate the bolt so I could reach the nut.  None of the bolts moved that I was able to reach.

Steve, we're not doing final final alignment today (we will do it tomorrow), so please go around and double-check my work by checking all of the dogs first thing in the morning.  Thanks.

 Almost all chamber dog clamps on the floor checked. There are a few exception where it is impossible to to get to the nut. 

Only the OOC nuts turned little bit. So our elastomer discs are holding up well. This means that the chamber anchoring to the floor is good.

  7326   Fri Aug 31 10:16:02 2012 janosch, SteveUpdate ETMX, scattering preps

The alignment of the pick-off mirror near ETMX is done. Everything turned out to be easy once we realized that there is no sense getting the alignment laser (going through viewport to pick-off to ITMX) back to ETMX. It is only necessary to hit ITMX somehow, since this makes sure that there is one scattered beam that will make it from ITMX to pick-off through viewport.

After the auxiliary optic (that we never used in the end) was removed again, we levelled the optical table.

So in the current setup, we can have small-angle scattering measurements on ITMX and large-angle scattering measurements on ETMX.

Attachment 1: IMG_1609.JPG
IMG_1609.JPG
Attachment 2: IMG_1608b.jpg
IMG_1608b.jpg
  7325   Fri Aug 31 07:32:49 2012 SteveUpdateSUStarget for BS

Quote:

We installed beam targets on PRM and BS suspension cages.

On both suspensions one of the screw holes for the target actually houses the set screw for the side OSEM.  This means that the screw on one side of the target only goes in partial way.

The target installed on BS is wrong!  It has a center hole, instead of two 45 deg holes.  I forgot to remove it, but it will obvious it's wrong to the next person who tries to use it.  I believe we're supposed to have a correct target for BS, Steve?

The earthquake stop screws on PRM were too short and were preventing installation of the PRM target.  Therefore, in order to install the target on PRM I had to replace the earthquake stops with ones Jenne and I found in the bake lab clean room that were longer, but have little springs instead of viton inserts at the ends.  This is ok for now, but

WE NEED TO REMEMBER TO REPLACE EARTHQUAKE STOPS ON PRM WHEN WE CLOSE UP.

We checked the beam through PRM and it's a little high to the right (as viewed from behind).  Tomorrow we're going to open ITMX chamber so that we can get a closer look at the spot on PR2.

 The two eye  target for  the BS is in the clean tool box. It actually has irises.

  7324   Thu Aug 30 20:35:09 2012 jamieUpdateSUStarget installed on PRM, temporary earthquake stops in place

We installed beam targets on PRM and BS suspension cages.

On both suspensions one of the screw holes for the target actually houses the set screw for the side OSEM.  This means that the screw on one side of the target only goes in partial way.

The target installed on BS is wrong!  It has a center hole, instead of two 45 deg holes.  I forgot to remove it, but it will obvious it's wrong to the next person who tries to use it.  I believe we're supposed to have a correct target for BS, Steve?

The earthquake stop screws on PRM were too short and were preventing installation of the PRM target.  Therefore, in order to install the target on PRM I had to replace the earthquake stops with ones Jenne and I found in the bake lab clean room that were longer, but have little springs instead of viton inserts at the ends.  This is ok for now, but

WE NEED TO REMEMBER TO REPLACE EARTHQUAKE STOPS ON PRM WHEN WE CLOSE UP.

We checked the beam through PRM and it's a little high to the right (as viewed from behind).  Tomorrow we're going to open ITMX chamber so that we can get a closer look at the spot on PR2.

  7323   Thu Aug 30 20:31:35 2012 JenneUpdateIOOPZT1 and PZT2 set to center of their ranges

[Koji, Jenne]

Jamie and Koji pointed out that we need to be doing the in-vac alignment with the PZTs at the center of their ranges.  Also, we confirmed that they were set to "closed loop off", so the strain gauges were not supplying any feedback.

PZT1 was set to 0 for both pitch and yaw, since it has a very limited range of motion right now, so 0 is close enough.

For PZT2, Koji and I moved the slider in pitch and yaw, and watched the LCD output monitor on the PZT driver at the bottom of 1Y3.  We saw the value on the LCD change between slider values +4 to -6 for PZT2 yaw, so it is set to -1 as the center.  We saw the value on the LCD change between slider values -4 to +5 for PZT2 pitch, so it is set to +0.5 as the center.   Beyond these slider values (the sliders all go -10 to +10), the LCD value didn't change, either at 0, or at the maximum. 

Since PZT1 doesn't really move, this shouldn't affect any of the alignment work that Suresh and I did last night, although we should quickly confirm tomorrow. On the agenda for tomorrow is adjusting PZT2 such that we hit the center of PR2 (and hopefully that will also put us through the center of the PRM target, if the alignment was done well enough last time), so it's okay that we have only now set it to the center of its range.

  7322   Thu Aug 30 20:20:52 2012 JenneUpdateSUSWatek camera placed on SE viewport of ITMX to look at PRM

[EricQ, Jenne]

We placed the Watek camera on the SE viewport of the ITMX chamber, and focused it on the face of PRM.  We are not able to see any scattered light transmitted through the PRM, so this camera was an ineffective way to try to check spot centering on the PRM.  Jamie placed one of the new targets on the PRM cage - see his elog for details.

To get more use of the camera, we need to mount it on something, at the 5.5 inch beam height, and then cover that something with clean foil so we can place the camera on the table, in the beamline in various places.  We also need to carefully wrap the cables in foil so the don't dirty anything inside.

  7321   Thu Aug 30 19:08:08 2012 JenneUpdateVACDogs on BS, ITM chambers checked

I tightened as many of the dog clamps on the bottom of the BS, ITMX and ITMY chambers as I could find.  I used a torque wrench at 45 ft-lbs.  Some of the bolts of the dogs were too long, and I couldn't find an extender thing to accommodate the bolt so I could reach the nut.  None of the bolts moved that I was able to reach.

Steve, we're not doing final final alignment today (we will do it tomorrow), so please go around and double-check my work by checking all of the dogs first thing in the morning.  Thanks.

  7320   Thu Aug 30 18:01:05 2012 ElliUpdateIOOSetting up Input MC cavity scan measurement

Riju, Elli

Today tried to take our first cavity scan.  We unplugged the 55MHz sideband input from the RF combiner on the PSL table, and connected a network analyser instead.  Using the network analyzer we injected a 12dBm signal (swept from 32MHz to 45MHz) through the RF combiner into the EOM to create our swept sidebands.  We measured the  MC cavity response by looking at the signal comming out of the RF photodiode on the MC2 table.  I replaced the BNC cable connected to the RF PD with a longer BNC cable that could reach our network analyzer next to the PSL table.  Riju will post a diagram of our setup.

We didn't see the expected carrier resonances when we performed a cavity scan.  The light incident on the RF PD is around 0.7micro Watts and we are still thinking about whether this is strong enough to see our signal above the noise.  We also want to work out what the strength of our swept sidebands is.  We will attempt to do a 'real' cavity scan tomorrow.

 

  7319   Thu Aug 30 17:03:32 2012 janosch, Manasa, SteveUpdate ETMX

The baffle has been moved away from ETMX towards the edge of the table (in fact, it is a little beyond the edge). It is also rotated so that its long edge is horizontal. In this way it was possible to center the baffle hole with respect to the optical axis, but also make it possible that the camera looks over the baffle.

We have tried to get an alignment beam from view port -> ETMX pick-off ->ITMX-> back to EX. This work was pretty much unsuccessful though. We could see the green laser scattering around ITMX, but there was no good way to know when the beam hit ITMX. So tomorrow we will find a better way to check where the beam is hitting at ITMX and finish the alignment of the scattering pick-off mirror.

  7318   Thu Aug 30 13:10:41 2012 jamieUpdateCamerasETMX

Quote:

We have done some work at ETMX today. We installed the baffle and placed two mirrors on the table.

The baffle position/orientation still needs to be checked more thoroughly to make sure that the beam will pass through the center of the baffle hole.

I must say that I am not at all happy with the baffle situation.  It is currently completely blocking our camera view of the ETMX face.  Here's a video capture of the ETMX face camera:

etmx-face-baffle.png

The circle is the baffle hole, through which we can see just the bottom edge of the test mass.  I don't think whatever benefit the baffle gives out weights the benefit of being able to see the spot on the mirror.

This afternoon we will try to adjust the baffle, and maybe the camera view mirror, to see if we can get a better shot of the center of the TM.  If we can see the beam spot through the hole we can probably live with it.  If not, I think we should remove the baffle.

  7317   Thu Aug 30 12:01:27 2012 janosch, Manasa,SteveUpdate ETMX

We have done some work at ETMX today. We installed the baffle and placed two mirrors on the table.

The baffle position/orientation still needs to be checked more thoroughly to make sure that the beam will pass through the center of the baffle hole.

One of the two mirrors will stay on the table as pickoff. The other is only temporarily installed for alignment purposes. Later today we will shoot a laser into the chamber that will reflect off one of these mirrors towards the center of ITMX, then go back to the pickoff mirror next to ETMX and hopefully make it through the viewport.

To place the pickoff mirror, we had to move the "cable rack" next to ETMX a few inches towards the back of the table.

  7316   Thu Aug 30 08:37:11 2012 SteveUpdateSUSETMX and ITMX needs more attention

Quote:

ETMX appears to be fine.  It was stuck to its OSEMs in the usual way.  I touched it and it dislodged and is now swinging freely.  Damping loops have been re-engaged.

Screenshot.png

Earthquake m4.1 test for ETMX and moves ITMX.  ITMX-LR sensor 0.3V

Attachment 1: after4.1mAtm.png
after4.1mAtm.png
Attachment 2: eq4.1mAtm.png
eq4.1mAtm.png
  7315   Thu Aug 30 08:12:39 2012 SteveUpdateGeneralIn vac plans for tomorrow, 8/30

 

 1,PRM spot can be viewed directly from the window south-east of ITMX chamber.  I can easy set up the mobile- Watek for this reason or you can just use an IR viewer.

   Remember, we have 2 SOS centering targets ready to use , that Rana was suggesting.

2, PR2 spot centering can be viewed directly through window north-west of ITMX.

3, We should put back the BS view pick-up mirror for the vertical camera on the BS chamber and adjust its upper pick-up.

4, The BS centering can be viewed with the mobile-Watek placed inside the BS chamber immediately.

  7314   Thu Aug 30 00:08:34 2012 jamieUpdateGeneralIn vac plans for tomorrow, 8/30

Quote:

We need to check spot centering on PRM with camera tomorrow.

Suresh checked that we're not clipped by IP ANG/POS pickoff mirrors, but we haven't done any alignment of IP ANG/POS.

 I think we should NOT do any adjustment of IP ANG/POS now.  We should in fact try to recover them when doing the PRM spot centering

Quote:

Tomorrow:  Open ITMX door.  Check with Watek that we're hitting center of PRM.  Then look to see if we're hitting center of PR2.  Then, continue through the chain of optics.

The motivation for removing the ITMX door was so that the scatter measurement team could check alignment of the new viewing mirror next to ETMX.  After discussion today we decided that everything can be done at the X end.  They can inject a probe beam into the ETMX chamber, bounce it off of ITMX and align the viewing mirror with the reflection.  So we'll leave ITMX door on for now.

We should, however, inspect the situation ITMY and make sure we have good clearance in the Y arm part of the Michaelson.  Koji previously expressed suspicion that we might have clipping on the southern edge of the POY steering mirror, so we need to check that out.

Koji and I discussed the situation for getting camera face views of BS and PRM.  Koji said the original idea was to see if we could install something at the south-east view port of ITMX chamber.  Steve also suggested utilizing the "ceiling" camera mounted on the top of the IOO chamber.

Vertex tasks:

  • check spot centering in PRM
  • check that REFL is getting cleanly to the AP table
  • check IPPOS and IPANG - we should be adjusting IPPOS or IPANG at this point
  • check spot centering on BS
  • remove ITMY north door
  • check clearance of POY steering mirror
  • ...

in parallel:

  • Steve will inspect the situation for getting a camera view of BS and PRM face, either through IOO or ITMX.

End X tasks:

  • install baffle
  • install "permanent" ITMX viewing mirror, on west side of ETMX - this might require moving ETMX SUS cable bracket south
  • install temporary steering mirror for probe laser on south-east side of ETMX
  • at some point the scatter guys can also do transmission measurements of the ETMX view ports
  • ...
  7313   Wed Aug 29 21:02:45 2012 JenneUpdateIOOOptics between Faraday and PRM are centered, realigned

[Jenne, Suresh, with support from Jamie and Koji]

MC spots measured, MC1, MC3 no change.

No clipping going through Faraday.

Beam hitting to the right of center of PZT1.  It was translated sideways so we are now hitting it on the center.  Knobs adjusted so we hit center of MMT1.

Beam totally obscured by Faraday on the way to MMT2.  MMT2 moved north, so that we clear the Faraday by more than a beam diameter.  MMT1 knobs adjusted to hit center of MMT2.

MMT2 knobs adjusted to hit center of PZT2.

PZT2 didn't have enough range with knobs, so we loosened it, pointed then adjusted with knobs so we're hitting center of PRM. 

We need to check spot centering on PRM with camera tomorrow.

Suresh checked that we're not clipped by IP ANG/POS pickoff mirrors, but we haven't done any alignment of IP ANG/POS.

 

Tomorrow:  Open ITMX door.  Check with Watek that we're hitting center of PRM.  Then look to see if we're hitting center of PR2.  Then, continue through the chain of optics.

  7312   Wed Aug 29 20:43:23 2012 KojiUpdateIOOSetup for a cavity scan or the input mode cleaner

The technique is based on detection of the beating between the resonant carrier and a resonant higher order mode.
This means that the beat signal is cancelled out if the transmitted beam is integrated over the entire beam.
Thus, you may want to introduce intentional clipping (or cutting a half of the beam with a razor blade). 

Ref: LIGO DCC G080467: Precise Measurements on Longitudinal and Transverse Mode Spacings of an Optical Cavity using an Auxiliary Laser

I am quite curious on the measurement as I am going to do the same measurement for the aLIGO OMC.
I am interested in seeing the statistical evaluation on the precision of the measurement.

  7311   Wed Aug 29 19:28:41 2012 Elli KingUpdateLSCSetup for a cavity scan or the input mode cleaner

 Riju, Elli

Today we prepared our experimental setup to take a cavity scan of the input mode cleaner, which we want to measure in the next day or so.  Attached is a diagram of our setup.

What we want to do is to inject a set of sidebands into the PSL and sweep their frequency from 32-45 MHz (a range just over one fsr of the mode cleaner- vfsr=11MHz).  We will measure the power transmitted out of the MC using a photo-diode and demodulate this signal with our input signal from the Marconi.  From this we should be able to see the resonant frequencies of the carrier and the higher order modes.

One aspect we spent some time thinking about; whether we would be able to inject a signal into an EOM given the EOM and the Marconi are not perfectly impedance matched.  Based on Kiwamu’s previous e-log entries designing the EOM, we decided that injecting a signal in 32-45 MHz region at 15dBm is similar to injecting the 29.5MHz sideband (at the same power level with very similar input impedance.) Fingers crossed we don’t blow anything up first week on the job.

Attachment 1: 40m_cavity_scan_diagram.jpg
40m_cavity_scan_diagram.jpg
  7310   Wed Aug 29 17:35:34 2012 KojiUpdatePSLPower reduced!

The biggest reason why we could not lock the MC was that the beam was not properly hitting the MC REFL diode.

Now the MC REFL DC is about ~0.1 and 1.2 when the MC is and is not locked.

We increased the power according to the quantitative analysis of the intracavity power in this earlier entry

Autolocker script for the low power MC was modified so that the initial VCO gain is 3 in stead of 10.
The 2 steps of super boost were also enabled again.

  7309   Wed Aug 29 17:09:57 2012 jamieUpdateSUSETMX OK, free swinging

ETMX appears to be fine.  It was stuck to its OSEMs in the usual way.  I touched it and it dislodged and is now swinging freely.  Damping loops have been re-engaged.

Screenshot.png

  7308   Wed Aug 29 17:02:41 2012 ericqUpdatePSLPower reduced!

Quote:

We have now reduced the power being input to the MC from 1.25W to 10mW, and changed out the MC refl BS for a mirror. 

The power was reduced via the PBS we introduced in Entry 7295.

While we were in there, we took a look at the AS beam, which was looking clipped on the monitor. Jenne felt that it appears that the clipping seems to be occurring inside the vacuum, possibly on the faraday. This will be investigated during the vent. 

 The power has been increased to 20mW. We got the 10mW number from the linked elog entry above. However, after venting we were having problems locking the MC. Upon investigating past elog posts, we found that 20mW was actually the power used in the past. The MC will now autolock. 

  7307   Wed Aug 29 12:46:49 2012 SteveUpdateVACVent completed in 4 hrs

 

ANTS ALERT please watch out for ants. We have them in the control room.

 

 

Attachment 1: vent72.png
vent72.png
Attachment 2: beforeVentOpl.png
beforeVentOpl.png
Attachment 3: beforeVent.png
beforeVent.png
  7306   Wed Aug 29 11:47:21 2012 ericqUpdateVACVenting

 [Steve, Eric]

I've been helping Steve vent this morning. The following things were done (from Steve's logbook):

  • Particle counts: 0.5 micron particles, 4200 counts per cubic ft
  • Vertex crane drive checked to be ok
  • Optical Levers set for local damping only
  • Saved some screens
  • PSL shutter and green shutters closed
  • HV Off checked, JAM nuts checked
  • Vac: Close V1, VM1, ans - VA6, open VM3 - RGA, cond: chamber open mode
  • 8AM: VV1 open to N2, regulator set  to 14 psi
  • 8:23AM: 35psi Instrument grade Air

(At this point, I took over the air canisters, while Steve made preparations around the lab. 

  • 9:00AM: 2nd air cylinder, 14 psi 
  • 9:40AM: 3rd air cyl
  • 10:20AM: 4th air cyl
  • 11:00AM: 5th air cyl

With the 5th cylinder, we began approaching 1 atm, so we slowed the regulator down to 5psi. Around 750 torr, Steve opened VV1 to air.

According to Steve, we will be at atmospheric pressure at  ~12:30pm.

  7305   Wed Aug 29 09:35:03 2012 janoschMetaphysicsRingdownripples 2

Ok, so the whole idea that mirror motion can explain the ripples is nonsense. At least, when you think off the ringdown with "pump off". The phase shifts that I tried to estimate from longitudinal and tilt mirror motion are defined against a non-existing reference. So I guess that I have to click on the link that Koji posted...

Just to mention, for the tilt phase shift (yes, there is one, but the exact expression has two more factors in the equation I posted), it does not matter, which mirror tilts. So even for a lower bound on the ripple time, my equation was incorrect. It should have the sum over all three initial tilt angles not only the two "shooting into the long arms" of the MC.

Quote:

Laser frequency shift = longitudinal motion of the mirrors

Ringing: http://www.opticsinfobase.org/ol/abstract.cfm?uri=ol-20-24-2463

Quote:

Hmm. I don't know what ringing really is. Ok, let's assume it has to do with the pump... I don't see how the pump laser could produce these ripples. They have large amplitudes and so I always suspected something happening to the intracavity field. Therefore I was looking for effects that would change resonance conditions of the intracavity field during ringdown. Tilt motion seemed to be one explanation to me, but it may be a bit too slow (not sure yet). Longitudinal mirror motion is certainly too slow. What else could there be?

 

 

  7304   Tue Aug 28 20:23:54 2012 KojiMetaphysicsRingdownripples

Laser frequency shift = longitudinal motion of the mirrors

Ringing: http://www.opticsinfobase.org/ol/abstract.cfm?uri=ol-20-24-2463

Quote:

Hmm. I don't know what ringing really is. Ok, let's assume it has to do with the pump... I don't see how the pump laser could produce these ripples. They have large amplitudes and so I always suspected something happening to the intracavity field. Therefore I was looking for effects that would change resonance conditions of the intracavity field during ringdown. Tilt motion seemed to be one explanation to me, but it may be a bit too slow (not sure yet). Longitudinal mirror motion is certainly too slow. What else could there be?

 

  7303   Tue Aug 28 19:21:37 2012 janoschMetaphysicsRingdownripples

Hmm. I don't know what ringing really is. Ok, let's assume it has to do with the pump... I don't see how the pump laser could produce these ripples. They have large amplitudes and so I always suspected something happening to the intracavity field. Therefore I was looking for effects that would change resonance conditions of the intracavity field during ringdown. Tilt motion seemed to be one explanation to me, but it may be a bit too slow (not sure yet). Longitudinal mirror motion is certainly too slow. What else could there be?

Quote:

Isn't it just a ringing of the intracavity power as you shifted the laser frequency abruptly?

Quote:

Let's see if the ripples observed in the MC ringdown can be due to tilt motion of the mirrors.

The time it takes to produce a phase shift corresponding to N multiples of 2*pi is given by:

t = sqrt(2*N*lambda/(L*omega_T^2*(alpha_1+alpha_2)))

L is the length of the MC (something like 13m), and alpha_1, alpha_2 are the DC tilt angles of the two mirrors "shooting into the long arms of the MC" produced by the MC control with respect to the mechanical equilibrium position. omega_T is the tilt eigenfrequency of the three mirrors (assumed to be identical). lambda = 1.064e-6m;

The time it takes from N=1 to N=2 (the first observable ripple) is given by: tau1 = 0.6/omega_T*sqrt(lambda/L/(alpha_1+alpha_2))

The time it takes from N=2 to N=3 is given by: tau2 = 0.77*tau1

etc

First, we also see in the measurement that later ripples are shorter than early ripples consistent with some accelerated effect. The predicted ripple durations tau seem to be a bit too high though. The measurements show something like a first 14us and a late 8us ripple. It depends somewhat on the initial tilt angles that I don't know really.

In any case, the short ripple times could also be explained if the tilt motions start a little earlier than the ringdown, or the tilt motion starts with some small initial velocity. The next step will be to program a little ringdown simulation that includes mirror tilts and see what kind of tilt motion would produce the ripples exactly as we observe them (or maybe tilt motion cannot produce ripples as observed).

 

 

ELOG V3.1.3-